SAT Math · Probability and conditional probability · medium

The probability of rain on any given day this week is 0.3, independent of other days. What is the probability of at least one rainy day…

The probability of rain on any given day this week is 0.3, independent of other days. What is the probability of at least one rainy day over the next 2 days?

Answer and explanation

Answer: C: 0.51

Use the complement. The chance of no rain on a single day is 1−0.3=0.71 - 0.3 = 0.7. The days are independent, so the chance of no rain on both days is 0.7×0.7=0.490.7 \times 0.7 = 0.49. The chance of at least one rainy day is 1−0.49=0.511 - 0.49 = 0.51, choice C. Choice A (0.6) adds the probabilities, 0.3+0.30.3 + 0.3, which double-counts the case where it rains both days. Choice B (0.09) is the probability of rain on both days, 0.3×0.30.3 \times 0.3. Choice D (0.3) is the probability for a single day and ignores the second day.

More Probability and conditional probability questions

All Probability and conditional probability questions

Learn the method behind questions like this one

The SAT Math course teaches each tested skill through lessons, step-by-step solutions to sample problems, the complete study guide and full-length practice tests.